# Invariants for $\mathbb{G}_{(r)}$-modules

### v1

@@ 1. Jordan types and specialization
Let $A$ be a Noetherian commutative $k$-algebra and consider a projective $A$-module $M_{A}$ of rank $m$ equipped with a $p$-nilpotent endomorphism $\theta_{A}$. As usual, we give $\operatorname{Spec}\nolimits A$ the Zariski topology. Give $\mathcal{Y}$ the topology whose closed subsets are those subsets $S\subset\mathcal{Y}$ satisfying the condition that if $s\in S$ and $s^{\prime}\leq s$ then $s^{\prime}\in S$.
Then
$JT_{A,M_{A}}(-):\operatorname{Spec}\nolimits A\quad\to\quad\mathcal{Y}_{\leq,m}$
@@
where $\overline{\{\underline{a}\}}$ is the closure of the singleton subset $\{\underline{a}\}$. Thus, it suffices to prove that $JT_{A,M_{A}}^{-1}(\overline{\{\underline{a}\}})$ is closed for all $\underline{a}\in\mathcal{Y}_{\leq,m}$ (since $\mathcal{Y}_{\leq,m}$ is finite). By definition, $JT_{A,M_{A}}^{-1}(\overline{\{\underline{a}\}})$ consists of those $x\in\operatorname{Spec}\nolimits A$ such that $JT_{A,M_{A}}(x)\ \leq\ \underline{a}$ in $\mathcal{Y}_{\leq,m}.$ By Proposition 1.3(4), $JT_{A,M_{A}}(x)\ \leq\ \underline{a}$ if and only if the rank of $(\theta_{x})^{s}:k(x)\otimes_{A}M_{A}\to k((x)\otimes_{A}M_{A}$ is less than or equal to the rank of $t^{s}$ on $\oplus_{i=1}^{p}(k[t]/t^{i})^{\oplus a_{i}}$ (which equals $\sum_{i=s+1}^{p}a_{i}\cdot(i-1-s)$) for all $s$, $1\leq s<p$.
@@ 2. Jordan type functions for $\mathbb{G}$-modules
For an affine group scheme $\mathbb{G}$ of finite type over $k$, we shall be interested in formulation invariants of finite dimensional $\mathbb{G}$-modules. We denote by $F^{r}:\mathbb{G}\to\mathbb{G}^{(r)}$ the $r$-th iterate of the Frobenius map $F:\mathbb{G}\to\mathbb{G}^{(1)}$, and we set
$\mathbb{G}_{(r)}\ \equiv\ ker\{F^{r}:\mathbb{G}\to\mathbb{G}^{(r)}\},$
an infinitesimal group scheme of height $\leq r$ with coordinate algebra $k[\mathbb{G}_{(r)}]$. Thus, $k[\mathbb{G}_{(r)}]$ is a finite dimensional Hopf algebra whose $k$-linear dual we denote by $k\mathbb{G}_{(r)}$.
For any commutative $k$-algebra $R$, we denote by $\mathbb{G}_{R}$ the group scheme $\mathbb{G}\times\operatorname{Spec}\nolimits R$ over $\operatorname{Spec}\nolimits R$ defined as the base change of the affine group scheme $\mathbb{G}$. Thus, the coordinate algebra of $\mathbb{G}_{R}$ equals $R\otimes k[\mathbb{G}]$. If $\mathbb{G}$ is a finite group scheme, then the “group algebra” of $\mathbb{G}_{R}$ (namely, $Hom_{R}(R\otimes k[\mathbb{G}],R)$ equals $R\otimes k\mathbb{G}$.
@@
Let $\mathbb{G}$ be an affine group scheme of finite type over $k$ and let $r$ be a positive integer. As proved in [22, Thm1.5], the functor of commutative $k$-algebras sending an algebra $A$ to the set of maps of group schemes $\mu:\mathbb{G}_{a(r),A}\to\mathbb{G}_{A}$ over $A$ (i.e., the set of height $r$ 1-parameter groups of $\mathbb{G}_{A}$) is representable by the spectrum $V_{r}(\mathbb{G})$ of the finitely generated commutative $k$-algebra $k[V_{r}(\mathbb{G})]$. Each such map $\mu$ factors through $\mathbb{G}_{(r),A}\hookrightarrow\mathbb{G}_{A}$.
Given a $\mathbb{G}_{(r)}$-module $M$ and a scheme theoretic point $x\in V_{r}(\mathbb{G})$, base change provides a $\mathbb{G}_{(r),k(x)}$-module structure on $k(x)\otimes M$. If $\mu_{x}:\mathbb{G}_{a(r),k(x)}\to\mathbb{G}_{(r),k(x)}$ is the 1-parameter subgroup over $k(x)$ parametrized by $\chi_{x}:\operatorname{Spec}\nolimits k(x)\to V_{r}(G)$, then restricting along $\mu_{x}$ determines a $\mathbb{G}_{a(r),k(x)}$-module structure on $k(x)\otimes M$.
In other words, the map of “group algebras” $\mu_{x,*}:k(x)\otimes k\mathbb{G}_{a(r)}\to k(x)\otimes k\mathbb{G}_{(r)}$ determines by restriction along $\mu_{x,*}$ the structure of a $k(x)\otimes k\mathbb{G}_{a(r)}$-module on $k(x)\otimes M$.
Remark 2.2.
A theorem of Suslin, Bendel, and the author ([23, Thm 5.2] asserts that there is a “canonical homomorphism $\psi:H^{ev}(\mathbb{G}_{(r)},k)\to k[V_{r}(\mathbb{G})]$ whose kernel is nilpotent and whose image contains every $p^{r}$-th power. This enables the authors to formulate in terms of 1-parameter subgroups a “rank variety interpretation” of cohomological support varieties for $\mathbb{G}_{(r)}$-modules.
For general infinitesimal group schemes we lack a concrete description of $V_{r}(\mathbb{G})$. In Section 4, we restrict our attentions to $\mathbb{G}$ of exponential type of height $r$ for which we do have such a description of $V_{r}(\mathbb{G})$.
The following definition makes explicit how the map of group algebras $\mu_{x,*}:k(x)\otimes k\mathbb{G}_{a(r)}\to k(x)\otimes k\mathbb{G}_{(r)}$ naturally determines by “restricting along $\epsilon_{r}:k[u]/u^{p}\to k\mathbb{G}_{a(r)}$” the map of $k(x)$-algebras $(\mu_{x})_{*}\circ\epsilon_{r}:k(x)[u]/u^{p}\to\ k(x)\otimes k\mathbb{G}_{(r)}$. We remind the reader that group algebra $k\mathbb{G}_{a(r)}$ (i.e., the dual of the coordinate algebra $k[\mathbb{G}_{a(r)}]=k[t]/t^{p^{r}}$) is naturally identified as the commutative $k$-algebra $k[u_{0},\ldots,u_{r-1}]/(u_{i}^{p})$ where $u_{i}$ is the linear dual to $t^{p^{i}}$.
@@
(2.3.1) $(\mu_{x})_{*}\circ\epsilon_{r}:k(x)[u]/u^{p}\to k(x)\otimes k\mathbb{G}_{a(r)}\to k(x)\otimes k\mathbb{G}_{(r)}\quad u\mapsto(\mu_{x})_{*}(u_{r-1}).$
Thus, if $M$ is a $\mathbb{G}$-module, then $((\mu_{x})_{*}\circ\epsilon_{r})^{*}(k(x)\otimes M)$ is equipped with a $k(x)[u]/u^{p}$-module structure such that $u$ acts as $(\mu_{x})_{*}(u_{r-1})$ on $k(x)\otimes M$.
To unpack $(\mu_{x})_{*}(u_{r-1})\in k(x)\otimes k\mathbb{G}_{(r)}$, we observe that this is the $k(x)$-linear function associated the the $k$-linear function $k[\mathbb{G}]\to k$ sending $f\in k[\mathbb{G}]$ to the coefficient of $t^{p^{r-1}}$ in the expansion of $\mu_{x}^{*}(f)\in k(x)\otimes k[\mathbb{G}_{a(r)}]=k(x)[t]/t^{p^{r}}$.
Maps of the form (2.3.1) are $\pi$-points of $\mathbb{G}$; $\pi$-point equivalence classes of these maps are in natural 1-1 correspondence with the scheme theoretic points of the $\Pi$-point scheme $\Pi(\mathbb{G})$ which is isomorphic to $\operatorname{Proj}\nolimits(-)$ applied to $H^{*}(\mathbb{G},k)$ (see [12]).
In the following definition, we recall the natural grading of $V_{r}(\mathbb{G})$ as discussed in [22, §1]. This is particularly relevant because $\mathbb{P}V_{r}(\mathbb{G})\equiv\operatorname{Proj}\nolimits k[V_{r}(\mathbb{G})]$ is naturally homeomorphic to the $\pi$-point scheme $\Pi(\mathbb{G})$ (see [15, Prop 3.1]).
@@
We consider $A$-points of $V_{r}(\mathbb{G}_{a(r)})$ as additive polynomials in $A[t]/t^{p^{r}}\to A[t]/t^{p^{r}}$ (i.e., maps of $k$ algebras given by $t\mapsto\sum_{i}a_{i}t^{p^{i}})$ and consider the monoid action $V_{r}(\mathbb{G}_{a(r)})\times\mathbb{A}^{1}\to V_{r}(\mathbb{G}_{a(r)})$ sending $(\sum_{i=0}^{r-1}a_{i}t^{p^{i}},\ b)$ to $\sum_{i=0}^{r-1}a_{i}(bt)^{p^{i}}.$
Let $\mathbb{G}$ be an affine group scheme of finite type over $k$. The right action of $\mathbb{A}^{1}$ on $V_{r}(\mathbb{G}_{a(r)})$ induces a right action of $\mathbb{A}^{1}$ on $V_{r}(\mathbb{G})$ with associated grading on $k[V_{r}(\mathbb{G})]$. We set
@@
If $\mu:\mathbb{G}_{(r),R}\to\mathbb{G}_{R}$ is an $R$-point of $V_{r}(\mathbb{G})$ for some commutative $k$-algebra $R$ and $s$ an $R$-point of $A^{1}$ (i.e., an element $s\in R$), then $(s\circ\mu)_{*}(u_{i})=s^{p^{i}}\mu_{*}(u_{i})$ (see the proof of [23, Prop 6.5]).
The following proposition is a rephrasing of Lemma 1.12 of [22].
@@
The graded commutative algebra $k[V_{r}(GLN)]$ is a graded quotient of the graded algebra $k\{X_{i,j}^{\ell};1\leq i,j,\leq N,0\leq\ell<r\}$ provided that $X_{i,j}^{\ell}$ is given degree $p^{\ell}$. Moreover, if $\mathbb{G}\subset GL_{N}$ is a closed immersion, then $k[V_{r}(\mathbb{G})]$ is a graded quotient of $k[V_{r}(GLN)]$.
[…]
@@
We identify the graded coordinate algebra of $V_{r}(\mathbb{G}_{a})=V_{r}(\mathbb{G}_{(r)})$ as the graded polynomial algebra $k[T_{0},T_{1},\ldots,T_{r-1}]$, where $T_{i}$ is given degree $p^{i}$. Then $\operatorname{Proj}\nolimits$ applied to the graded algebra is the weighted projective space $w\mathbb{P}(1,p,\ldots,p^{r-1})$. Let $k[t_{0},t_{1},\ldots,t_{r-1}]$ denote the graded polynomial algebra with each $t_{i}$ of degree 1. The graded map of $k$-algebras
@@
determine a ramified covering map
@@
[…]
Observe that $U_{2}\ \subset\ w\mathbb{P}(1,p,p^{2})$ has an isolated singularity at $T_{0}=0=T_{1}$; the maximal ideal at this singular point is generated by the $p+1$ elements of
$\{T_{0}^{p^{2}}/T_{2},T_{0}^{p^{2}-p}T_{1}/T_{2},\ldots,T_{0}^{p}T_{1}^{p-1}/T_{2},T_{1}^{p}/T_{2}\}$.
The representability by $k[V_{r}(\mathbb{G})]$ of the functor sending a commutative $k$-algebra $A$ to set of all height $r$ 1-parameter subgroups $\mathbb{G}_{a(r),A}\to\mathbb{G}_{A}$ (over $\operatorname{Spec}\nolimits A$) leads to the universal 1-parameter subgroup $\mathcal{U}_{\mathbb{G},r}$ as recalled in the following definition.
@@
([14, §2]) Let $\mathbb{G}$ be an affine group scheme of finite type over $k$ and let $r$ be a chosen positive integer. Then the universal 1-parameter subgroup over $V_{r}(\mathbb{G})$
(2.8.1) $\mathcal{U}_{\mathbb{G},r}:\mathbb{G}_{a(r),k[V_{r}(\mathbb{G})]}\ \to\mathbb{G}_{(r),k[V_{r}(\mathbb{G})]}$
of height $r$ is the 1-parameter subgroup of the group scheme $\mathbb{G}_{(r),k[V_{r}(\mathbb{G})]}$ over $V_{r}(\mathbb{G})$ corresponding to the identity map $V_{r}(\mathbb{G})\to V_{r}(\mathbb{G})$. By the universality of $\mathcal{U}_{\mathbb{G},r}$, for any scheme-theoretic point $x\in V_{r}(\mathbb{G})$ the restriction along $\chi_{x}:\operatorname{Spec}\nolimits k(x)\to V_{r}(G)$ of $\mathcal{U}_{\mathbb{G},r}$ equals the map $\mu_{x}:\mathbb{G}_{a(r),k(x)}\ \to\mathbb{G}_{(r),k(x)}$ of Definition 2.1.

### v2

@@ 1. Jordan types and specialization
Let $A$ be a Noetherian commutative $k$-algebra and consider a projective $A$-module $M_{A}$ of rank $m$ equipped with a $p$-nilpotent endomorphism $\theta_{A}$. As usual, we give $\operatorname{Spec}\nolimits A$ the Zariski topology. Then
$JT_{A,M_{A}}(-):\operatorname{Spec}\nolimits A\quad\to\quad\mathcal{Y}_{\leq,m}\quad\hookrightarrow\quad\mathcal{Y}_{\leq}$
@@
where $\overline{\{\underline{a}\}}$ is the closure of the singleton subset $\{\underline{a}\}$. Thus, it suffices to prove that $JT_{A,M_{A}}^{-1}(\overline{\{\underline{a}\}})$ is closed for all $\underline{a}\in\mathcal{Y}_{\leq,m}$ (since $\mathcal{Y}_{\leq,m}$ is finite). By definition, $JT_{A,M_{A}}^{-1}(\overline{\{\underline{a}\}})$ consists of those $x\in\operatorname{Spec}\nolimits A$ such that $JT_{A,M_{A}}(x)\ \leq\ \underline{a}$ in $\mathcal{Y}_{\leq,m}.$ By Proposition 1.3 (4), $JT_{A,M_{A}}(x)\ \leq\ \underline{a}$ if and only if the rank of $(\theta_{x})^{s}:k(x)\otimes_{A}M_{A}\to k((x)\otimes_{A}M_{A}$ is less than or equal to the rank of $t^{s}$ on $\oplus_{i=1}^{p}(k[t]/t^{i})^{\oplus a_{i}}$ (which equals $\sum_{i=s+1}^{p}a_{i}\cdot(i-s)$) for all $s$, $1\leq s<p$.
@@
Notation 1.14.
In the remainder of this paper, we shall leave implict the subscript $\leq$ on $\mathcal{Y}_{\leq,m}$ and $\mathcal{Y}_{\leq}$.
@@ 2. Jordan type functions for $\mathbb{G}$-modules
Consider an affine group scheme $\mathbb{G}$ of finite type over $k$. We are interested in formulating invariants for finite dimensional $\mathbb{G}_{(r)}$-modules. Here, $\mathbb{G}_{(r)}$ is the kernel of the $r$-th iterate of the Frobenius map $F:\mathbb{G}\to\mathbb{G}^{(1)}$, where $k[\mathbb{G}^{(1)}]$ is the base change of $k[\mathbb{G}]$ along the $p$-th power map $(-)^{p}:k\to k$. (See [18, §1].) Thus,
$\mathbb{G}_{(r)}\ \equiv\ ker\{F^{r}:\mathbb{G}\to\mathbb{G}^{(r)}\}$
is an infinitesimal group scheme of height $\leq r$ with coordinate algebra $k[\mathbb{G}_{(r)}]$. We denote by $k\mathbb{G}_{(r)}$ the $k$-linear dual of the finite dimensional Hopf algebra $k[\mathbb{G}_{(r)}]$.
For any commutative $k$-algebra $R$, we shall employ various notations (depending upon context) for
$\mathbb{G}_{R}\quad\equiv\quad\mathbb{G}\times\operatorname{Spec}\nolimits R\quad\equiv\quad R\otimes\mathbb{G}$
for the group scheme defined as the base change along $k\to R$ of the affine group scheme $\mathbb{G}$ over $k$. Thus, the coordinate algebra of $\mathbb{G}_{R}$ equals $R\otimes k[\mathbb{G}]$. If $\mathbb{G}$ is a finite group scheme over $k$, then the “group algebra” of $\mathbb{G}_{R}$ (namely, $Hom_{R}(R\otimes k[\mathbb{G}],R))$ is identified with $R\otimes k\mathbb{G}$ by sending $1\otimes\phi\in R\otimes k\mathbb{G}$ to $id_{R}\cdot\phi(-):R\otimes k[\mathbb{G}]\to R$.
@@
Let $\mathbb{G}$ be an affine group scheme of finite type over $k$ and let $r$ be a positive integer. As proved in [26, Thm1.5], the functor of commutative $k$-algebras sending an algebra $A$ to the set of maps of group schemes $\mu:\mathbb{G}_{a(r),A}\to\mathbb{G}_{A}$ over $A$ (i.e., the set of height $r$ 1-parameter subgroups of $\mathbb{G}_{A}$) is representable. The commutative $k$-algebra representing this functor is denoted $k[V_{r}(\mathbb{G})]$, its spectrum by $V_{r}(\mathbb{G})$.
Given a $\mathbb{G}_{(r)}$-module $M$ and a scheme theoretic point $x\in V_{r}(\mathbb{G})$, base change provides a $\mathbb{G}_{(r),k(x)}$-module structure on $k(x)\otimes M$. Namely, if $\mu_{x}:\mathbb{G}_{a(r),k(x)}\to\mathbb{G}_{(r),k(x)}$ is the 1-parameter subgroup over $k(x)$ parametrized by $\chi_{x}:\operatorname{Spec}\nolimits k(x)\to V_{r}(G)$, then restricting along $\mu_{x}$ determines a $\mathbb{G}_{a(r),k(x)}$-module structure on $k(x)\otimes M$.
In other words, restriction along the map of “group algebras” $\mu_{x,*}:k(x)\otimes k\mathbb{G}_{a(r)}\to k(x)\otimes k\mathbb{G}_{(r)}$ determines the $k(x)\otimes k\mathbb{G}_{a(r)}$-module structure on $k(x)\otimes M$.
The following definition makes explicit how the map of group algebras $\mu_{x,*}:k(x)\otimes k\mathbb{G}_{a(r)}\to k(x)\otimes k\mathbb{G}_{(r)}$ naturally determines by “restricting along $\epsilon_{r}:k[u]/u^{p}\to k\mathbb{G}_{a(r)}$” a map of $k(x)$-algebras $(\mu_{x})_{*}\circ\epsilon_{r}:k(x)\otimes k[u]/u^{p}\to\ k(x)\otimes k\mathbb{G}_{(r)}$. We remind the reader that the group algebra $k\mathbb{G}_{a(r)}$ (i.e., the dual of the coordinate algebra $k[\mathbb{G}_{a(r)}]=k[t]/t^{p^{r}}$) can be described as the commutative $k$-algebra $k[u_{0},\ldots,u_{r-1}]/(u_{i}^{p})$ where $u_{i}$ is the $k$-linear dual to $t^{p^{i}}$.
@@
(2.2.1) $(\mu_{x})_{*}\circ\epsilon_{r}:k(x)\otimes k[u]/u^{p}\to k(x)\otimes k\mathbb{G}_{a(r)}\to k(x)\otimes k\mathbb{G}_{(r)},\ u\mapsto(\mu_{x})_{*}(1\otimes u_{r-1}).$
Thus, if $M$ is a $\mathbb{G}$-module, then $(\mu_{x})_{*}\circ\epsilon_{r})^{*}(k(x)\otimes M)$ is equipped with a $k(x)\otimes k[u]/u^{p}$-module structure such that $u$ acts as $(\mu_{x})_{*}(1\otimes u_{r-1})$ on $k(x)\otimes M$.
To unpack $(\mu_{x})_{*}(u_{r-1})\in k(x)\otimes k\mathbb{G}_{(r)}$, we observe that this is the $k(x)$-linear function sending $f\in k[\mathbb{G}]$ to the coefficient of $t^{p^{r-1}}$ in the expansion of $\mu_{x}^{*}(f)\in k(x)\otimes k[\mathbb{G}_{a(r)}]=k(x)[t]/t^{p^{r}}$.
We recall the local Jordan type of a finite dimensional $\mathbb{G}_{(r)}$-module as first formulated in [15] and [16] by Pevtsova and the author.
@@
(See [16, Cor 3.8].) Let $\mathbb{G}$ be an affine group scheme of finite type over $k$, $r\geq 1$, and $M$ a finite dimensional $\mathbb{G}_{(r)}$-module. Then the local Jordan type of $M$ at a scheme-theoretic point $x\in V_{r}(\mathbb{G})$, $JT_{\mathbb{G},r,M}(x)$, is defined to be the Jordan type of the $k(x)[u]/u^{p}$-module $((\mu_{x})_{*}\circ\epsilon_{r})^{*}(k(x)\otimes M)$.
In the following definition, we recall the natural grading on $V_{r}(\mathbb{G})$ as discussed in [26, §1]. This is particularly relevant because there is a natural morphism $\operatorname{Proj}\nolimits k[V_{r}(\mathbb{G})]\to\mathbb{P}H^{*}(\mathbb{G}_{(r)},k)\simeq\Pi(\mathbb{G}_{(r)})$ which is a homeomorphism of Zariski spaces to the $\pi$-point scheme $\Pi(\mathbb{G})$ (see [17, Prop 3.1]).
Definition 2.4.
The natural right action $V_{r}(\mathbb{G}_{a})\times\mathbb{A}^{1}\to V_{r}(\mathbb{G}_{a})$ sends
$(\nu:\mathbb{G}_{a(r),A}\to\mathbb{G}_{a(r),A},\alpha\in A)$ to $\alpha\cdot\nu:\mathbb{G}_{a(r),A}\to\mathbb{G}_{a(r),A}$ defined by $(\alpha\cdot\nu)^{*}(t)=\nu^{*}(\alpha\cdot t)$. Composing this with 1-parameter subgroups $\mu:\mathbb{G}_{a(r),A}\to\mathbb{G}_{(r),A}$ defines a right action $V_{r}(\mathbb{G})\times\mathbb{A}^{1}\to V_{r}(\mathbb{G})$. This action corresponds to a grading on the commutative algebra $k[V_{r}(\mathbb{G})]$.
We set
@@
If $\mu:\mathbb{G}_{(r),R}\to\mathbb{G}_{R}$ is an $R$-point of $V_{r}(\mathbb{G})$ for some commutative $k$-algebra $R$ and $s$ an $R$-point of $A^{1}$ (i.e., an element $s\in R$), then $(s\circ\mu)_{*}(u_{j})=s^{p^{i}}\mu_{*}(u_{j})$ for $u_{j}\in k\mathbb{G}_{a(r)}=k[u_{0},\cdots,u_{r-1}]/(u_{i}^{p}).$ (See the proof of [27, Prop 6.5]).)
We rephrase Lemma 1.12 of [26] in order to elaborate upon this grading of $V_{r}(\mathbb{G})$.
@@
The commutative algebra $k[V_{r}(GL_{N})]$ is a graded quotient of the symmetric algebra over $k$ generated with generators $\{X_{i,j}^{\ell};1\leq i,j,\leq N,0\leq\ell<r\}$ provided that $X_{i,j}^{\ell}$ is given degree $p^{\ell}$. Moreover, if $\mathbb{G}\subset GL_{N}$ is a closed immersion, then $k[V_{r}(\mathbb{G})]$ is a graded quotient of $k[V_{r}(GL_{N})]$.
Consequently, for any affine group scheme $\mathbb{G}$ of finite type over $k$, $k[V_{r}(\mathbb{G})]$ is a graded commutative algebra generated by homogeneous elements of degrees $p^{i},\ 0\leq i<r$.
The following theorem summarizes fundamental results of Suslin, Bendel, and the author which relate 1-parameter subgroups of $\mathbb{G}$ of height $r$ to cohomological support varieties of $\mathbb{G}_{(r)}$.
Theorem 2.6.
[26, Thm 5.2], [27, Thm 6.7], [27, Cor 6.8] Let $\mathbb{G}$ be an affine group scheme of finite type over $k$, $r$ a positive integer, and $M$ a finite dimensional $\mathbb{G}_{(r)}$-module. There is a natural morphism of commutative graded $k$-algebras $\psi:H^{ev}(\mathbb{G}_{(r)},k)\ \to\ k[V_{r}(\mathbb{G})]$ of degree $\frac{p^{r}}{2}$ with associated morphism $\Psi:V_{r}(\mathbb{G})\ \to\ \operatorname{Spec}\nolimits H^{ev}(\mathbb{G}_{(r)},k)$ of schemes which is a homeomorphism on scheme-theoretic points with the Zariski topology. This morphism restricts to a homeomorphism of closed subspaces
$\Psi:V_{r}(\mathbb{G})_{M}\quad\stackrel{{\scriptstyle\approx}}{{\to}}\quad\operatorname{Spec}\nolimits H^{ev}(\mathbb{G}_{(r)},k)_{M},$
where
$V_{r}(\mathbb{G})_{M}\quad=\quad\{x\in V_{r}(\mathbb{G}):\ J_{\mathbb{G},r,M}(x)\not=\frac{dimM}{p}\cdot[p]\}$
and $\operatorname{Spec}\nolimits H^{ev}(\mathbb{G}_{(r)},k)_{M}$ is the (affine) cohomological support variety of $M$.
Moreover, $\psi$ determines the “projectivized morphism”
$\Psi:\operatorname{Proj}\nolimits V_{r}(\mathbb{G})\ \to\ \operatorname{Proj}\nolimits H^{ev}(\mathbb{G}_{(r)},k)$ which restricts to a homeomorphism
$\Psi:\operatorname{Proj}\nolimits V_{r}(\mathbb{G})_{M}\ \stackrel{{\scriptstyle\approx}}{{\to}}\ \operatorname{Proj}\nolimits H^{ev}(\mathbb{G}_{(r)},k)_{M}.$
Remark 2.7.
We refer the reader to [23] for some explorations of $V_{r}(\mathbb{G})$, including the reducibility and lack of equi-dimensionality in general, and some bounds on dimensions.
Remark 2.8.
The map $(\mu_{x})_{*}\circ\epsilon_{r}$ of (2.2.1) is a $\pi$-point of $\mathbb{G}_{(r)}$ whose equivalence class $(\mu_{x})_{*}\circ\epsilon_{r}]$ is a scheme-theoretic point of the $\pi$-point space $\Pi(\mathbb{G}_{(r)})$. This determines the map of points induced by the homeomorphism
(2.8.1) $\mathbb{P}V_{r}(\mathbb{G})\quad\stackrel{{\scriptstyle\approx}}{{\to}}\quad\mathbb{P}H^{*}(\mathbb{G}_{(r)},k)\quad\stackrel{{\scriptstyle\sim}}{{\to}}\quad\Pi(\mathbb{G}_{(r)})$
obtained by composing $\Psi$ of Theorem 2.6 with the natural isomorphism $\mathbb{P}H^{*}(\mathbb{G}_{(r)},k)\stackrel{{\scriptstyle\sim}}{{\to}}\Pi(\mathbb{G}_{(r)})$ of [14, Thm 7.5].
Observe that sending the 1-parameter subgroup $\mu_{x}:\mathbb{G}_{a(r),k(x)}\to\mathbb{G}_{k(x)}$ to the equivalence class of the $\pi$-point $(\mu_{x})_{*}\circ\epsilon_{r}:k(x)\otimes k[u]/u^{p}\to k(x)\otimes k\mathbb{G}_{a(r)}\to k(x)\otimes k\mathbb{G}$ determines a natural projection
(2.8.2) $V_{r}(\mathbb{G}){\text{-}}\{0\}\quad\to\quad\mathbb{P}V_{r}(\mathbb{G}).$
For any finite dimensional $\mathbb{G}_{(r)}$-module $M$, the maps of (2.8.1) restrict to homeomorphisms of corresponding support varieties for $M$
$\mathbb{P}V_{r}(\mathbb{G})_{M}\ \stackrel{{\scriptstyle\approx}}{{\to}}\ \mathbb{P}H^{*}(\mathbb{G}_{(r)},k)_{M}\ \stackrel{{\scriptstyle\sim}}{{\to}}\ \Pi(\mathbb{G}_{(r)})_{M}.$
Remark 2.9.
The more general theory of $\pi$-points supports developed by Pevtsova and the author in [14], $M\to\Pi(H)_{M}$, applies to $H$-modules for an arbitrary finite group scheme $H$. A disadvantage of considering $M\to\Pi(H)_{M}$ rather than $M\mapsto V_{r}(\mathbb{G})_{M}$ is that points of $\Pi(H)$ are equivalence classes of certain flat maps of $K$-algebras of the form $\alpha_{K}:K[u]/u^{p}\to KH$ for field extensions $K/k$. For maximal or generic points of $\Pi(H)$, the Jordan type of a finite dimensional $H$-module $M$ is independent of the choice of representative of such an an equivalence class $\alpha_{K}]$, but for other points of $\Pi(H)$ the Jordan type of $M$ depends upon a choice of representative of the equivalence class. See [17]).
@@
We identify the graded coordinate algebra of $V_{r}(\mathbb{G}_{a})=V_{r}(\mathbb{G}_{a(r)})$ with the graded polynomial algebra $k[T_{0},T_{1},\ldots,T_{r-1}]$, where $T_{i}$ is given degree $p^{i}$. Then $\operatorname{Proj}\nolimits$ applied to the graded algebra is the weighted projective space $w\mathbb{P}(1,p,\ldots,p^{r-1})$. Let $k[t_{0},t_{1},\ldots,t_{r-1}]$ denote the graded polynomial algebra with each $t_{i}$ of degree 1. The graded map of $k$-algebras
@@
determines a ramified covering map
@@
$U_{2}\ =\ \operatorname{Spec}\nolimits k[T_{0}^{ip}T_{1}^{p-i}/T_{2};\ 0\leq i\leq p].$
Observe that $U_{2}\ \subset\ w\mathbb{P}(1,p,p^{2})$ has an isolated singularity at $T_{0}=0=T_{1}$; the maximal ideal at this singular point is generated by the $p+1$ elements
$T_{0}^{p^{2}}/T_{2},T_{0}^{p^{2}-p}T_{1}/T_{2},\ldots,T_{0}^{p}T_{1}^{p-1}/T_{2},T_{1}^{p}/T_{2}$.
The representability by $k[V_{r}(\mathbb{G})]$ of the functor sending a commutative $k$-algebra AA to the set of all height $r$ 1-parameter subgroups $\mathbb{G}_{a(r),A}\to\mathbb{G}_{A}$ (over $\operatorname{Spec}\nolimits A$) leads to the universal 1-parameter subgroup $\mathcal{U}_{\mathbb{G},r}$ as recalled in the following definition.
@@
([16, §2]) Let $\mathbb{G}$ be an affine group scheme of finite type over $k$ and let $r$ be a chosen positive integer. Then the universal 1-parameter subgroup of height $r$ for $\mathbb{G}$,
($2.1$) $\mathcal{U}_{\mathbb{G},r}:\mathbb{G}_{a(r),k[V_{r}(\mathbb{G})]}\ \to\mathbb{G}_{(r),k[V_{r}(\mathbb{G})]},$
is the 1-parameter subgroup for the group scheme $\mathbb{G}_{(r),k[V_{r}(\mathbb{G})]}$ over $k[V_{r}(\mathbb{G})]$ corresponding to the identity map $V_{r}(\mathbb{G})\to V_{r}(\mathbb{G})$. By the universality of $\mathcal{U}_{\mathbb{G},r}$, for any scheme-theoretic point $x\in V_{r}(\mathbb{G})$ the restriction along $\chi_{x}:\operatorname{Spec}\nolimits k(x)\to V_{r}(\mathbb{G})$ of $\mathcal{U}_{\mathbb{G},r}$ equals the map $\mu_{x}:\mathbb{G}_{a(r),k(x)}\ \to\mathbb{G}_{(r),k(x)}$ of Definition 2.1.